### v8

# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

Abstract
Current AI systems demonstrate remarkable capabilities yet remain specialised, in part because no unified measure of general intelligence has been established. Existing evaluation frameworks, which focus primarily on language or perception tasks, offer limited insight into generality. The Artificial General Intelligence Testbed (AGITB) introduces a complementary benchmarking suite of fourteen elementary tests, with thirteen implemented as fully automated procedures.
AGITB evaluates models on their ability to forecast the next input in a temporal sequence, step by step, without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants, such as determinism, sensitivity, and generalisation, that parallel principles of biological information processing. Designed to resist brute-force or memorisation-based strategies, AGITB enforces unbiased and autonomous learning. The human cortex satisfies all tests, whereas no current AI system meets the full AGITB criteria, demonstrating its value as a rigorous, interpretable, and actionable benchmark for evaluating progress toward artificial general intelligence. A reference implementation of AGITB is freely available on GitHub.
Keywords: artificial general intelligence, benchmarking, generalisation, symbol grounding problem, temporal sequence prediction
1 Introduction
Despite great achievements in machine learning and neural network architectures, artificial intelligence (AI) systems still lack the flexibility and robustness of human intelligence. Marcus2020 correctly observed that although large language models (LLMs) can generate highly fluent outputs, they rely primarily on statistical pattern recognition rather than grounded, compositional reasoning. Mitchell2025 further summarised why even the latest large reasoning models (LRMs), despite some impressive achievements, cannot be trusted. Surface-level competence obscures a deeper lack of understanding, which is an essential prerequisite for artificial general intelligence (AGI).
The progress in AI has led to a growing speculation that AGI is near. However, such claims remain difficult to substantiate without a rigorous and informative metric. Assessing progress toward AGI requires more than specialised metrics or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
Various attempts have been made to define such tools, the most iconic being the Turing test proposed by Turing1950. However, no existing test fully achieves its intended purpose; moreover, all lack key properties such as gradual resolution, interpretability, and full automatizability. The metrics typically assess surface-level competence rather than underlying mechanisms of generalisation. In response to these limitations, this paper introduces the Artificial General Intelligence Testbed (AGITB), a novel benchmark for validating foundational cognitive abilities in artificial intelligence systems.
AGITB draws inspiration from the Ladder to Human-Comparable Intelligence introduced by Sprogar2018, extending its conceptual foundations into a unified testbed comprising fourteen core requirements, with thirteen implemented as fully automated and domain-agnostic tests. Although AGITB does not aim to evaluate consciousness or semantic comprehension, it provides a principled framework for distinguishing narrow AI systems from those exhibiting generalisable, adaptive behaviour. To position AGITB within the broader landscape of AGI evaluation, we include a comparative analysis with the Abstraction and Reasoning Corpus (ARC; Chollet2019) and the NeuroBench framework (Yik2025).
2 Background
The rapid progress of deep learning has enabled AI systems to exhibit increasingly sophisticated reasoning, problem-solving, and dialogue capabilities. However, despite these advances, a persistent reluctance remains to attribute "intelligence" to machines. This hesitation is partly rooted in the intuitive association of intelligence with uniquely human traits, such as consciousness, self-awareness, and subjective experience, which remain elusive in artificial systems.
Historically, as AI systems have succeeded in domains once considered hallmarks of human intelligence, definitions of intelligence have undergone significant shifts. For instance, the success of Deep Blue against Garry Kasparov in chess (a task formerly seen as a benchmark for AGI) was quickly reframed as a triumph of brute-force computation rather than genuine intelligence. Such redefinitions risk obscuring real milestones. As AI approaches human-level capabilities, we may inadvertently set a perpetually receding goalpost for AGI, failing to recognise it even when achieved.
Although AGI is typically envisioned as matching human cognitive flexibility across diverse domains, its evaluation has largely defaulted to narrow, task-specific metrics. This is partly due to the absence of a universally accepted AGI benchmark. Researchers have thus gravitated toward achieving superhuman performance in discrete domains, where progress can be clearly quantified. However, such specialised benchmarks favour narrow AI by rewarding depth within isolated subdomains rather than breadth of adaptation and general reasoning—hallmarks of general intelligence. Ironically, some of these benchmarks are now so specialised that humans have difficulty with them.
2.1 A Benchmark That Only Humans and AGI Can Meet
An effective AGI test must be trivial to solve for humans yet remain inaccessible to contemporary machine learning models that rely on brute-force methods, pretraining, or statistical pattern matching. Such a test must demand capabilities that transcend memorisation or domain-specific heuristics, requiring generalisation, abstraction, and adaptive reasoning.
A valid AGI evaluation must either (1) reveal and exploit a fundamental cognitive gap between humans and machines or (2) define a behavioural capacity that current artificial systems cannot replicate. The first strategy is increasingly fragile, as advanced AI systems often bypass genuine understanding through massive pretraining on diverse datasets. The second strategy may involve a more faithful emulation of human cortical computation, pointing to neuromorphic architectures such as spiking neural networks, which more closely mirror the brain’s time-sensitive, event-driven dynamics, as described by Maass1997.
In alignment with this biologically grounded perspective, the AGITB departs from symbolic, high-level evaluations and instead assesses intelligence at the lowest, signal-processing level. While Turing was right to suggest that communication could serve as a basis for evaluating machine intelligence, natural language remains problematic as a test medium: it depends on shared human experiences and symbols whose meanings are ungrounded in machines, as argued by Harnad1990.
AGITB thus adopts a more elemental approach. Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalise patterns in raw binary signals. A neural spike by itself contains the smallest amount of information possible and is, as such, grounded but free of other semantics. A binary signal accurately represents the neural spike.
Building on the view of Hawkins2004 that intelligence is fundamentally about extracting structure from data to enable prediction, AGITB operates at the level of signal-based forecasting. This approach aligns closely with the functioning of biological intelligence at the cortical level, which processes time-sensitive sensory spike trains rather than disembodied symbols.
3 Artificial General Intelligence Testbed
The testbed supports the development and evaluation of AGI by defining a clear set of requirements that a model under evaluation must meet to qualify as generally intelligent. A model must satisfy all requirements in order to claim success on the benchmark.
The central premise of AGITB is that it represents a necessary (though not sufficient) condition for the existence of artificial general intelligence. In the absence of a clear definition of intelligence, this premise cannot be formally proven, though it could be falsified by the existence of a non-AGI system that satisfies all AGITB criteria. The assumption of necessity follows from the fact that, apart from the human brain, no known artificial system has yet fulfilled its requirements, suggesting that a solution lies beyond the current state of the art in artificial intelligence.
AGITB is not, in itself, a sufficient criterion for AGI; a system that passes all tests is not automatically a fully-fledged AGI, as it is not required to exhibit higher-level capabilities such as reasoning or natural language fluency.
3.1 Architecture
AGITB evaluates an AGI model as a black box that predicts the next input based on the historical sequence of observed signals. Each input consists of ten bits, the specific semantics of which are immaterial; each bit may represent an arbitrary channel, such as a pixel, an audio band, or actuator feedback.
Each input represents a snapshot of multiple parallel signals at a single time step (Figure 1a). Spatial organisation within each input encodes local structure, whereas semantic richness arises from the temporal evolution of the input sequence (Figure 1b). The interaction between spatial and temporal dimensions gives rise to structured patterns that are challenging for the model to adapt to.
Figure 1: (a) Example of a 10-bit input $x$ with four bits set.
(b) Example of an input sequence.
3.2 Operation
The testbed presents the AGI model with a stream of inputs over time. At each time step $t$, the model receives an input $x_{t}$ and is required to predict the subsequent input $x_{t+1}$, as shown in Figure 2. The central challenge is not simply extrapolation but discerning the underlying causes or regularities that produce the observed input stream and using that understanding to make accurate future predictions.
Figure 2: Iterative adaptation in discrete time. At the previous time step (t1t-1), the model issued the prediction $x_{t}^{*}$. After observing the realised input $x_{t}$, it adapted its internal state in response to the error in the second bit and subsequently produced the one-step-ahead prediction $x_{t+1}^{*}$.
3.2.1 Design
A distinguishing architectural feature of AGITB is its deliberate avoidance of conventional correctness metrics such as accuracy or mean-squared error. The limitations of these metrics are twofold. First, they cannot reliably distinguish between AGI and human performance, or that of non-AGI systems, as contemporary AI models can already surpass humans on standard benchmarks. Second, in an era of elevated expectations driven by specialised generative AI systems, the anticipated AGI performance on such metrics is often set unrealistically high. Even well-educated humans possessing fully developed and highly parallelised brains may underperform relative to current AI systems. It is therefore unreasonable to expect a first-generation AGI, potentially operating on a simplified and computationally constrained simulation of the brain, to match or exceed human-level results. Currently, our understanding of neural mechanisms is insufficient to justify such expectations.
Without conventional metrics or predefined performance thresholds, it can be difficult to determine whether a model has achieved a meaningful level of competence. AGITB addresses this challenge by employing a self-referential evaluation approach, in which the model under test is compared against its own independent instances. Each test constructs a controlled scenario involving one or more such models, whose internal states and behaviours are analysed comparatively. Requiring each test to be passed 100 times renders AGITB an extreme form of stress testing, ensuring that successful performance reflects genuine robustness rather than chance. Success is thus defined in terms of the relative consistency or superiority of model responses, rather than by any external quantitative metric.
For these reasons, and to minimise type I errors, AGITB employs an all-or-nothing criterion: the system under evaluation must successfully pass all fourteen tests. This design choice is justified, as individual tests are solvable by non-AGI systems, whereas the simultaneous satisfaction of all fourteen requirements is likely to demand capabilities beyond narrow intelligence. AGITB therefore posits that what non-AGI systems lack is the property referred to as "intelligence," understood as the synergistic integration of all fourteen tests.
3.3 The 14 AGI Requirements
The reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
Requirement 1 — Unbiased start
Assertion: A model starts unbiased.
Assertion: An unbiased model produces an empty prediction.
AGITB rests on the foundational assumption that general-purpose learning systems, such as the brain, do not begin with an innate understanding of external inputs but instead acquire meaning through interaction with their environment. Each system must construct semantic content from raw sensory data rather than rely on pre-encoded [^1] Although certain reflexes may be genetically specified, they do not constitute genuine understanding. As Hawkins2004 argue, low-level reflexes are not prerequisites for intelligence but rather evolutionary features of the subcortical “old brain”. knowledge. Accordingly, a model should originate in an unbiased initial state shaped exclusively by subsequent inputs, where “unbiased” denotes the absence of informational content rather than the lack of structural capacity for learning.
A model that produces non-empty predictions prior to receiving any input is considered biased, as it encodes unwarranted assumptions about the future. To ensure neutrality, models must initialise in a completely blank state and remain so until conditioned on actual data. Although a spike-free prediction could be interpreted as a bias toward silence, this convention is justified by biological analogy: neurons do not activate in the absence of input. Two blank model instances should be regarded as equivalent, as neither has been influenced by prior experience.
The unbiased start requirement serves as the foundation for other tests that operationalise the emergence of meaning in the absence of prior grounding.
Requirement 2 — Bias
Assertion: A model that has received input can no longer be regarded as unbiased.
Every input modifies the model’s state, shaping its internal dynamics. As such, the mere act of input processing introduces bias based on experience.
Requirement 3 — Determinism
Assertion: If two models are equivalent, they must have received the same inputs.
Biological neurons operate in a functionally deterministic manner, ensuring stability and consistency in brain function. Although minor stochastic effects may occur, they do not undermine the rule-governed nature of neural processing. By analogy, AGITB assumes that two functionally identical models must have experienced identical input histories: any difference in input necessarily produces a divergence in state. This principle reinforces the view that a model’s internal state is fully determined by its input history.
Determinism at the level of neural signal processing is necessary for stable, reproducible brain function, whereas the apparent unpredictability of cognition stems from the system’s complexity rather than from genuine indeterminacy (Cave2016).
Requirement 4 — Sensitivity
Assertion: Two distinct models remain distinct, even after prolonged exposure to identical inputs.
Cortical systems exhibit chaotic sensitivity to initial conditions: small variations in early experiences can lead to divergent trajectories over time. This deterministic sensitivity, amplified by complex internal interactions, underlies the apparent unpredictability of decision-making.
Requirement 5 — Time
Assertion: Changing the order of inputs results in a different model state.
Since the model’s state depends on its cumulative history, the sequence of inputs critically shapes learning and adaptation. The capacity to recognise and exploit temporal structure is therefore a defining property of intelligent systems.
However, sensitivity to input order alone is not sufficient. Because temporal coding makes time an intrinsic feature of neural data, Requirement 6 is designed to exclude models that perform equally well on temporally invariant data, as such models fail to make functional use of temporal information.
Requirement 6 — Refractory period
Assertion: The model must be able to adapt to any minimal-period input sequence that respects biologically plausible refractory periods.
Assertion: The model cannot adapt to an input sequence that repeats a neural spike in violation of refractory-period constraints.
Biological intelligence relies on discrete spikes for communication and learning. AGITB enforces refractory periods to reflect this constraint, embodying the biological principle that a neuron cannot fire again immediately after activation.
While refractory periods are not the source of spiking variability, they impose a minimum separation between spikes, preventing continuous or unmodulated firing. This constraint helps preserve the temporal diversity needed for synaptic adaptation. Gerstner2002 showed that input sequences that lack sufficient variability, such as constant or overly repetitive spiking, fail to support effective learning. Consequently, AGITB permits only those temporal sequences that respect biologically plausible refractory dynamics while remaining agnostic to any particular semantic encoding of signals.
This requirement does not exclude the possibility of more general solutions. A universal system might simulate compliance with this constraint without relying on genuine refractory dynamics, whereas a system based on refractory processing cannot its absence.
Requirement 7 — Temporal adaptability
Assertion: The model must be able to adapt to a temporal pattern with the user-specified period.
Assertion: The model must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
This test requires the model to learn and track temporal structure across multiple timescales. Unlike rigid pattern-matching systems, an intelligent model should detect and predict a recurring structure regardless of its exact periodicity. Systems that can accommodate only a single, predetermined temporal scale therefore fail this requirement.
Requirement 8 — Stagnation
Assertion: There exists a limit beyond which the model can no longer adapt, even to patterns that would otherwise be learnable.
Cognitive systems inevitably reach adaptation limits as their finite resources become saturated. This test evaluates whether such a boundary arises when the model is exposed over time to input sequences that would otherwise be learnable.
Requirement 9 — Content sensitivity
Assertion: Adaptation time depends on the content of the input sequence.
The structural complexity of an input sequence influences the rate at which a model adapts, where adaptation time is defined as the number of iterations required for the model to predict the entire temporal pattern accurately. Simple or highly regular sequences typically lead to rapid convergence, whereas irregular or noisy inputs demand longer exposure before the model can reliably capture and reproduce the underlying pattern.
Requirement 10 — Context sensitivity
Assertion: Adaptation time depends on the state of the model.
The model’s current state reflects the cumulative influence of past inputs and therefore provides the context in which new information is processed. When subsequent inputs are consistent with structure already established through prior learning, adaptation proceeds quickly. Conversely, when new inputs conflict with this learned context, the model requires additional time to reorganise its state before accurate prediction becomes possible.
Requirement 11 — Unobservability
Assertion: Distinct models may exhibit the same observable behaviour.
Identical external outputs may arise from distinct internal states. This many-to-one mapping highlights that observable behaviour alone cannot disclose the underlying structure or history of a model’s internal dynamics.
Requirement 12 — Denoising
Assertion: A model can recall a sequence despite perturbations.
An intelligent model should be able to recall a previously observed sequence, even when the inputs are perturbed by noise. When re-exposed to familiar stimuli, such a model is expected, on average, to outperform random guessing in predictive tasks. Average performance over the 20 runs is used as the evaluation criterion because random models may occasionally generate correct predictions by chance, without demonstrating genuine learning or structural understanding.
Test 12 is intentionally designed as a conservative benchmark. The model must outperform random guessing in all 100 independent trials. This procedure is not interpreted as a conventional significance test (e.g., at the 5% level); rather, it functions as a stringent robustness check. The design ensures that only large, systematic performance gains yield a passing result. Modest or marginal improvements, while potentially real, are intentionally regarded as failures, as the objective is to identify only clear and substantial advances in model capability.
Requirement 13 — Generalization
Assertion: The model performs above chance on previously unseen inputs.
Only models capable of generalisation can derive lasting benefit from prior learning. After exposure to a given set of stimuli, such models are expected, on average, to outperform random baselines when predicting future inputs. As in the preceding requirement, average performance over the 20 runs is used as the evaluation criterion because random models may occasionally generate correct predictions by chance, without demonstrating genuine learning or structural understanding.
The generalisation assessment follows the same conservative design as Test 12. By requiring success across all repetitions, the procedure imposes a stringent threshold in which a single underperforming trial results in failure. As a consequence, the test tolerates virtually no noise or marginal effects. This criterion is intended not to maximise statistical power but to enforce robustness, ensuring that only models exhibiting a clear and consistent advantage obtain a passing result.
Requirement 14 — Bounded Prediction Latency
A model must provide low-latency, signal-level predictions to ensure functional viability. Biological brains achieve approximately constant reaction times through extensive parallel processing, keeping per-input latency bounded and stable under expected workloads.
No universal automated test running on serial hardware can feasibly verify this property across all possible model implementations. Accordingly, it is the responsibility of model developers to ensure that their architectures support efficient parallel execution with bounded latency. In this respect, artificial neural networks, with their inherently parallelisable structure, offer a practical advantage.
3.4 Search space
To prevent models from relying on brute-force memorisation, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints in both time and memory. In AGITB, tasks typically involve predicting a random temporal sequence of seven (or more) binary inputs, each comprising ten bits. This yields a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
AGITB incorporates a biologically inspired refractory period, which prohibits any neuron (bit) from firing in consecutive time steps. This restriction substantially reduces the number of valid sequences. There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits under the condition that a $1$ never carries over to the next time step, where $F_{i}$ denotes the $i$-th Fibonacci number with $F_{0}=0$.
In some cases, AGITB further constrains the space by requiring the sequence to be cyclic, such that the first input also satisfies the refractory condition relative to the last input in the sequence. The number of distinct cyclic temporal sequences respecting the refractory constraint is $|S^{\prime\prime}|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
The choice of seven-step sequences with ten-bit inputs is sufficient to detect non-AGI behaviour while maintaining computational efficiency. Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing runtime. The current configuration therefore ensures that each test remains both computationally feasible and diagnostically informative.
Within the comparatively constrained AGITB environment, every randomly generated input sequence is, in principle, learnable through exposure. However, the sheer size of the search space makes any form of explicit teaching-to-the-test computationally infeasible. Given that real-world sensory inputs may ultimately encompass tens of thousands of bits, a genuine AGI system must employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from high-dimensional data.
3.5 Performance
3.6 Performance
Before assessing the utility of AGITB, it is essential to recognise its role as a pragmatic benchmark rather than an end in itself. Similar to the Turing Test, which Harnad1992 characterised as an empirical criterion rather than a philosophical manoeuvre, AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence. The ultimate objective remains the development of AGI, not merely achieving success on the benchmark.
AGITB yields meaningful insights only when developers adhere strictly to its core requirements. Misinterpretations of fundamental elements, such as the notion of an “empty” initial state, may lead to erroneous conclusions and impede genuine progress toward AGI.
Overall, AGITB provides a structured framework for empirically assessing the AGI potential of diverse computational paradigms, including classical symbolic architectures, artificial neural networks, and large language models. Before benchmarking artificial systems, however, it is necessary to consider the baseline performance of human cognition.
3.6.1 Human Performance
The inability to directly compare internal cortical states makes it impossible to verify AGITB requirements in humans in a strict computational sense. Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with basic cognitive competencies, it is reasonable to assume that humans naturally satisfy most requirements. Demand 1 (Unbiased start), however, warrants further discussion.
Owing to prior experience and cognitive bias, an adult human’s cortex may appear to “fail” the unbiased start prerequisite, as it is no longer in an unconditioned state and may generate non-empty predictions. AGITB, however, requires that the unbiased state precede the first input—a condition satisfied only in the fetal cortex. At that stage, the cortex lacks synaptic organisation and, upon receiving its first sensory input, meets the criterion of true neutrality.
Several of the more complex AGITB tests have cognitive-level analogues that can be observed through reasoning and introspection. Temporal flexibility (Requirement 7), for example, poses no difficulty for humans, who readily recognise temporal patterns of varying durations. Stagnation (Requirement 8) reflects the finite capacity of the human cortex to store and maintain knowledge; its behavioural analogue resembles the onset of cognitive saturation or early dementia, in which recent experiences are lost. Because humans learn different types of content at varying rates, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is immediately evident. The unobservability test (Requirement 11) formalises the fact that internal human mental states cannot be directly inspected.
The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when confronted with new or distorted inputs. Finally, bounded prediction latency (Requirement 14) is a well-established property of the human brain, which maintains approximately constant reaction times through extensive parallel processing.
3.6.2 Classical Symbolic Programs
In principle, two complementary designs for AGI can be conceived. One embeds, either explicitly or implicitly, expectations about the external world within the system. The other is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
The former category includes most AI and purported AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990. Although such systems may display behaviour that appears intelligent, their interpretations of symbols depend on programmer-supplied conventions rather than grounded understanding, and they therefore cannot qualify as genuine AGI.
More specifically, systems that embed expectations inevitably incorporate the designer’s assumptions about the meaning of the signals they process. In classical symbolic architectures, the program itself constitutes prior knowledge: its rules and representations presuppose interpretations of the symbols being manipulated. The very existence of such a program violates AGITB’s first test, which prohibits external knowledge of any kind. In effect, the AGI program smuggles in the symbol-grounding problem it is meant to avoid.
AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary signals. Only an expectation-free system could, in principle, satisfy this condition. This perspective suggests that a genuine AGI may not explicitly encode the operations of intelligence, but rather the dynamics of a substrate from which intelligence can emerge—a view aligned with the “Brain Simulator Reply” proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980). To date, however, no such system is known.
3.6.3 Artificial Neural Network Performance
A similar dichotomy applies to connectionist architectures: ANNs can, in principle, be constructed either with built-in expectations (introduced through pretraining regimes or architectural priors) or as expectation-free systems that begin from a neutral initial state.
AGITB’s requirement that models begin in a completely unbiased, knowledge-free state stands in fundamental tension with the dominant paradigm of modern deep learning. Contemporary neural models typically rely on extensive pretraining, during which network weights are shaped by prior exposure to structured or labelled data. Moreover, by mapping symbolic inputs to numerical vectors, standard ANNs effectively shift the symbol-grounding problem into a number-grounding problem. Although these vector representations capture relational regularities within the training data, they also introduce spurious associations not anchored in real-world semantics, leading to the phenomenon commonly described as hallucination.
Only an expectation-free network can be regarded as unbiased. Yet a network without any initialised weights exhibits no effective learning dynamics: its units are not halted, but remain dormant and produce no informative output. This highlights a fundamental limitation of current ANN architectures: they do not initiate learning autonomously but instead depend on an external training procedure to drive adaptation. AGITB, by contrast, requires a blank system capable of initiating adaptation autonomously in an unfamiliar environment. To date, no such mechanism has been demonstrated in artificial neural networks.
3.6.4 Large Language Model Performance
Because a large language model is a type of artificial neural network, it fails the unbiased start test (Requirement 1). LLMs cannot learn entirely from scratch, as their behaviour is shaped in advance by the linguistic patterns and data on which they were pretrained. Although an LLM’s internal state is readily accessible and, unlike that of a human, can be inspected or compared across instances, any such evaluations remain moot until the first requirement is satisfied.
Two possibilities warrant consideration. The first is whether LLMs, while not constituting AGI, might nevertheless simulate one. For example, an LLM could be prompted to imitate a model receiving binary inputs as specified by AGITB. However, before examining such simulations, it is necessary to consider the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs. This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window. Once this window is exceeded, earlier tokens are discarded or compressed, weakening or eliminating the long-term dependencies required by the test.
The second possibility is whether an LLM can autonomously generate a solution to AGITB when provided with the reference implementation. In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purported to satisfy the requirements, none progressed beyond the adaptability test (Requirement 7).
In summary, LLMs do not perform genuine learning solely from prompts, nor can they acquire the grounded, context-dependent understanding characteristic of human cognition. These limitations extend to large reasoning models, which share the same underlying architectural foundations.
3.7 Remarks
AGITB evaluates a model’s predictive capabilities after exposure to temporal sequences of both structured and random inputs. Random input sequences with arbitrary internal correlations are employed to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream itself rather than from prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals; the random inputs need not resemble real-world sensory data.
The low-level, binary operational framework makes AGITB particularly well suited for evaluating NeuroAI models that aim to satisfy principles of the embodied Turing Test, as proposed by Zador2023, wherein cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to higher-level abstraction mirrors the broader trajectory of AI, from early perceptrons to large-scale models such as GPT.
Hand-engineered, task-specific systems have historically struggled to scale toward genuine generalisation, underscoring the need for adaptive, learning-based architectures such as artificial neural networks. However, state-of-the-art neural systems rely heavily on pretraining with structured or symbolic data, a process that inevitably introduces biases and bypasses the grounding of meaning in raw sensory input. This raises a central and unresolved question: how can such models initiate learning directly from unstructured input, as required to satisfy AGITB?
3.8 Cheating the Benchmark
Because AGITB’s tests are individually simple to solve, one might imagine circumventing the benchmark by engineering task-specific solutions and having the model selectively deploy them depending on the detected test scenario. In principle, the task being administered could be inferred by monitoring the number of instantiated models and the sequence of invoked methods.
However, such an approach would amount to subverting the benchmark rather than advancing AGI research. Although AGITB could be hardened against this form of cheating (by, for example, shuffling tests or redesigning the programming interface) these measures would reduce the transparency and interpretability of the testbed and thereby hinder its intended use by human developers.
For this reason, the AGITB reference implementation is kept deliberately readable and accessible. To date, no artificial system has demonstrated the level of performance that AGITB requires. Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing a more obfuscated version of the testbed.
4 Competing Benchmarks
Among existing benchmark tasks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation. A related effort is NeuroBench, which is designed to support the systematic evaluation of neuromorphic and other biologically inspired architectures. Both ARC and NeuroBench rely on a variety of correctness and complexity metrics to compare non-AGI models; their primary purpose is to distinguish weaker from stronger narrow systems. In contrast, AGITB is constructed to assess whether a model satisfies a set of necessary conditions associated with artificial general intelligence, rather than to rank models within a specific performance spectrum.
4.1 ARC
ARC presents visual reasoning tasks in which a model must infer novel transformations (such as recolouring, rearranging, or modifying spatial patterns) from a sequence of two input–output examples defined on discrete spatial grids.
However, ARC implicitly assumes the presence of high-level cognitive priors, including object permanence, spatial reasoning, numerical abstraction, and causal inference. These priors are not formally specified, placing an ambiguous and open-ended burden on the model designer. In contrast, AGITB adopts a fundamentally different stance: it treats the system under evaluation as a blank slate that must acquire structure and function exclusively through interaction with temporally structured input.
Although ARC presumes temporal reasoning, it does not explicitly test it, since each task presents only two images to illustrate a transformation. AGITB, by contrast, evaluates cognition as a dynamic process unfolding over time. A model can acquire knowledge and predictive capability only through continuous exposure to temporally structured data, not from disconnected before–after snapshots that lack the temporal continuity needed to infer causal relationships. For example, to recognise an object moving left, a model in AGITB must observe multiple intermediate states across time; the final image alone is insufficient to infer the transformation. Temporal structure, rather than static pattern comparison, provides the substrate from which invariants and causal relations can be learned.
In summary, ARC evaluates high-level intelligence grounded in human cognitive priors, whereas AGITB evaluates adherence to fourteen low-level computational principles intended to support the emergence of such priors. ARC and the Turing Test both frame intelligence through an anthropocentric lens, embedding assumptions drawn from human cognition. AGITB instead conceptualises intelligence as a universal capacity for learning that does not rely on innate symbolic structures or species-specific expectations.
4.2 NeuroBench
NeuroBench provides a unified framework for benchmarking diverse AI models across a standardised set of tasks and metrics. It is particularly oriented toward neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability. Within its algorithm track, the framework evaluates models on several challenges relevant to general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
The predictive modelling challenge, which involves forecasting chaotic functions, is most closely aligned with AGITB’s central premise that intelligence fundamentally concerns the prediction of future states. NeuroBench employs a synthetic one-dimensional Mackey-Glass time series for this task, a dataset designed for architectures with limited input/output capacity.
However, several issues limit the usability of chaotic function prediction (CFP) as a general AGI benchmark task. First, the Mackey-Glass data are numerical, and NeuroBench does not prescribe the encoding scheme. An inappropriate encoding can distort the temporal and causal structure of the observed signals, such that a numeric value—much like a symbol—derives its meaning from human interpretation rather than from the model’s own grounded understanding. This effectively reintroduces the symbol-grounding problem in a numerical form.
Second, the threshold for AGI-level performance is not clearly defined. Although symmetric mean absolute percentage error (sMAPE) is a standard forecasting metric, NeuroBench does not specify what performance level would correspond to general intelligence. Notably, humans themselves perform poorly at anomaly detection and long-horizon prediction of the Mackey-Glass signal (Thill2020).
Third, although long-term prediction is not inherently problematic, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to detect and correct its own errors. This design is incompatible with online learning, in which an AGI should continuously update its internal state upon observing discrepancies between prediction and outcome. NeuroBench, by contrast, emphasises offline learning and assumes that an AGI would behave as a purely mechanistic predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, and genuine agency.
4.3 Direct comparison
Table 1 highlights the key differences among the tasks used in the three benchmarks. Whereas ARC and NeuroBench presuppose or require models to exhibit high-level cognitive capacities (such as object recognition, spatial manipulation, and various forms of reasoning), AGITB instead focuses on minimal, precisely defined requirements that can be evaluated directly at the signal-processing level.
Property ARC CFP AGITB
Interface modality Visual Numeric Binary
AGI type Human Universal Universal
Cognitive priors Yes No No
Abstraction level High Medium Low
Formal AGI requirements No (qualitative) Partial Yes (explicit, testable)
Task preparation Manual Automatic Automatic
Symbol Grounding Problem Yes Yes No
Input dimensionality $200\times 100$ (4 colours) 1 number 10 bits
Temporal sequence length 2 750+ 7+
Table 1: Core properties of ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.
5 Conclusion
Although AGITB is proposed as a general testbed and benchmark for artificial general intelligence, it is nonetheless grounded in empirical knowledge of neural function in the human brain. This grounding is motivated by the fact that neuron-based systems—most notably biological brains and, to a more limited extent, artificial neural networks—are the only systems known to support a broad range of flexible cognitive abilities. By contrast, there is currently no evidence that fundamentally different computational architectures can give rise to comparable forms of general intelligence. Pursuing more abstract or architecture-agnostic criteria may therefore be premature until the principles underlying natural, neuron-based intelligence are better understood.
Unlike conventional benchmarks that focus on high-level task performance, such as question answering or language translation, AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex. It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework comprising fourteen fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
Unlike conventional benchmarks that target high-level task performance, such as question answering or language translation, AGITB evaluates whether a system exhibits behaviours associated with core operational principles of the biological cortex. Its focus is on low-level, biologically grounded computational properties that are believed to underlie the emergence of general intelligence. The testbed comprises fourteen tightly interdependent tests, each simple in isolation but collectively requiring the kind of adaptive, general-purpose learning expected of an AGI.
AGITB requires models to begin in an unbiased initial state and to acquire all functionality solely through exposure to structured or random input. This mirrors neuroscientific evidence that cortical learning is fundamentally input driven: neural circuits develop through experience, not through pre-encoded semantics. In biological systems, high-level cognition arises not from symbolic manipulation but from the continual adaptive prediction of low-level sensory signals. Such prediction is more than pattern matching; it supports the progressive construction of signal-grounded knowledge from which abstraction and generalisation can emerge.
AGITB is solvable by humans yet remains unsolved by classical algorithms and current state-of-the-art AI systems. This persistent performance gap provides empirical support for the claim that AGITB targets capabilities characteristic of general rather than narrow intelligence. Although the absence of an artificial solution does not constitute a formal proof of adequacy, the fact that humans succeed where machines do not indicates that the benchmark captures functionally relevant aspects of general intelligence. In this sense, AGITB serves as a discriminative test and a principled tool for steering the development of systems capable of genuinely general, adaptive learning.
Acknowledgments and Disclosure of Funding
The author acknowledges the financial support from the Slovenian Research Agency (research core funding No. P2-0057).
Appendix A LLM prompt
You are managing a binary-pattern prediction model participating in a signal-level AGI test. You will receive a 10-bit binary string (e.g., 0101010011) that represents a 10-bit sensory input to the model.
Your task is to simulate a model that **predicts the next 10-bit input**. The model has an internal state that only updates after receiving an input. Predictions must always be exactly 10 bits long, and the internal state can retain information from hundreds of past inputs.
Rules:
  1. 1.
    Upon receiving a 10-bit binary input (e.g., 1100010001):
    • Update the internal state of the model.
    • Generate a 10-bit prediction based on the state of the model.
    • Output the model’s 10-bit prediction.
  2. 2.
    **State updates occur only after input is received.**.
  3. 3.
    **Correct predictions are critical**:
    • Incorrect predictions trigger adaptation, updating the model’s internal state to improve future predictions.
    • The model should learn and adapt continuously.
  4. 4.
    Response format:
    • Each response should consist of a single 10-bit binary string on its own line.
    • Example: 0010101110
  5. 5.
    Formatting: Keep responses clean and minimal, with no explanations, commentary, or punctuation.
  6. 6.
    Begin by outputting the model’s prediction.
We begin now.